Crew 215 Science Report 19Nov2019

[title Science Report – November 19th]

Science Report 19 November 2019
SOL: 9
Crew 215 – Expedition Boomerang Crew
Submitted by Science Officer Andrew Wheeler
Science Dome Operations: The experimenting with gypsum and sand to create a form of
plaster of paris cement commenced. Various samples of known mass were heated for 5 or
10 minutes and either combined with sand or not and then mixed with 4ml or 8ml of
water. Each slurry was placed into a tray and is being allowed to dry to examine
strength. A byproduct of this experiment is the realization that the finely ground
gypsum is actual plaster of paris and could be used for its intended medical purpose on
Green Hab: Green Hab operations have continued with additional soil and potting mix
allowing further crops to be sown. See Green Hab report.
RAM Operations: N/A
EVA: The final EVA for science saw us returning rock specimens used for identification
and training purposes to the collection sites to the west of White Rock Canyon. See EVA
EVA Suit Maintenance: Nominal. See Operations Report.
Psycho-social Stressors Monitoring: Data collection continues through the hexoskin(TM)
monitoring vests.
Additional activities: Two potential medical uses of material being liberated from the
gypsum experiments, the original medical use of plaster of paris and the liberated
water (being equivalent to distilled water) used to create a saline solution (NaCl),
were evaluated concurrently with the cement creation trials.

Crew 215 Sol Summary 19Nov2019

[title Sol Summary – November 19th]

[category Sol-Summary ]

Crew 215 Sol Summary Report 19-NOV-2019
Sol: 9
Summary Title: The last EVA
Author’s name: Andrew Wheeler
Mission Status: Active
Sol Activity Summary: Gypsum experimentation got underway with the first samples being heated for either 5 or 10 minutes, either mixed with sand or not, reconstituted with 4ml or 8ml of water and set to dry in trays. During these experiments it became clear that the liberated water (distilled water equivalent) could be used to create a saline solution for medical needs and that the fine grained gypsum was actual plaster of paris and could be used for its original medical purpose. The morning also saw the last EVA as we returned training rock specimens to their original collection locations west of White Rock Canyon along 1101 Road (519500E 4247300N). The afternoon saw further experimentation with the gypsum samples and potting of new seedlings in the Green Hab. The sol was rounded out with Taco Tuesday, and fruit jelly with chocolate pudding for desert.
Look Ahead Plan: More plantings in the Green Hab and a continuation of the gypsum cement investigations.
Anomalies in work: N/A.
Weather: Possible rain overnight and during the day, light winds.
Crew Physical Status: Healthy
EVA: Sample return. See EVA Report.
Reports to be file: Sol Summary Report, Operations Report, Science Report, EVA Report, Green Hab Report, Journalist Report, Daily Photos.
Support Requested: N/A

Copyright © The Mars Society. All rights reserved. | Main Site